# Pulse Shaping Filter Design for Integrated Sensing & Communication with Zak-OTFS

### v1

@@ Abstract
Zak-OTFS is an emerging framework for integrated sensing & communication (ISAC) in high delay and Doppler spread environments. A critical enabler for ISAC with Zak-OTFS is the design of pulse shaping filters. For sensing, a localized pulse shaping filter enables ideal input-output (I/O) relation estimates close to the physical scattering channel. For communication, orthogonality of the pulse shape on the information lattice prevents inter-symbol interference, and no time and bandwidth expansion enables full spectral efficiency. A filter simultaneously meeting all three objectives is ideal for ISAC. Existing filter designs achieve two of the above objectives, but not all three simultaneously. For instance, the sinc filter is orthogonal and bandwidth/time-limited, but is not localized. The Gaussian filter is localized and bandwidth/time-limited, but not orthogonal. The RRC filter is localized and orthogonal, but not bandwidth/time-limited. A recently proposed hybrid Gaussian-sinc filter is more localized than the sinc filter and bandwidth/time-limited, but is not orthogonal. In this work, we design optimal pulse shaping filters meeting all three objectives via the Isotropic Orthogonal Transform Algorithm. The proposed pulse shaping filters offer improved data detection (communication) and I/O relation estimation (sensing) performance compared to existing filter choices in the literature.
@@ I Introduction
Zak-OTFS (orthogonal time frequency space modulation) [1, 2, 3] is an emerging framework for communications and radar sensing in doubly selective channels with high delay and Doppler spreads. The Zak-OTFS carrier waveform is a pulse in the delay-Doppler (DD) domain, formally a quasi-periodic localized function termed pulsone with specific periods along delay and Doppler. When the pulsone’s delay period exceeds the channel delay spread and the pulsone’s Doppler period exceeds the channel Doppler spread, the Zak-OTFS Input/Output (I/O) relation changes slowly and predictably, and does not fade [1, 2, 3]. This non-fading behavior of Zak-OTFS makes it well-suited for high rate communications in extreme mobility scenarios, such as on bullet trains, where legacy time-frequency modulations, such as orthogonal frequency division multiplexing (OFDM), fail. Moreover, the response to a single Zak-OTFS carrier provides an image of the scattering environment, which can be used to predict the I/O response to all other carriers. The image of the scattering environment supports sensing in systems such as passive radars that integrate sensing and communications [2].
In its original form, the DD pulsone occupies infinite time and bandwidth. For practical implementation, the DD pulsone is limited to a finite time interval $T$ and bandwidth $B$ via DD domain filtering with a pulse shaping filter [1, 2, 3, 4, 5, 6, 7, 8]. Subsequently, $BT$ information symbols are mounted on DD pulsones shifted by multiples of the delay resolution $\nicefrac{{1}}{{B}}$ and Doppler resolution $\nicefrac{{1}}{{T}}$. An ideal pulse shaping filter must satisfy three objectives: (i) localization to enable accurate I/O relation estimation (sensing), (ii) orthogonality on the information lattice to prevent inter-symbol interference, and (iii) no time and bandwidth expansion beyond $B$ and $T$ to achieve full spectral efficiency. A filter simultaneously meeting all three objectives is ideal for both sensing and communication.
@@
Previous works on Zak-OTFS utilize standard analytical pulse shaping filters, such as sinc [1, 2, 3, 4, 5], root raised cosine (RRC) [1, 2, 3, 4, 5, 6, 7] and Gaussian [5], which meet two but not all three objectives simultaneously. For instance, the sinc filter has the advantage of being orthogonal to shifts by multiples of the delay and Doppler resolution where information symbols are mounted. This property results in excellent data detection (communication) performance assuming the I/O relation is known perfectly. In the absence of perfect I/O relation knowledge, however, the high sidelobes of the sinc filter result in high I/O relation estimation (sensing) error, thus degrading practical data detection performance [1, 2, 3, 4, 5]. The RRC filter overcomes the high sidelobe drawback of the sinc filter at the cost of bandwidth and time expansion beyond $B$ and $T$ [1, 2, 3, 4, 5, 6, 7], making it less suitable for practical implementation. The Gaussian filter has the advantage of being maximally localized in the DD domain – enabling excellent I/O relation estimation (sensing) – however, is not orthogonal to shifts by multiples of the delay and Doppler resolution [5] and thus has poor data detection performance.
Recent work [8] utilizes the (normalized) product of the Gaussian and sinc filters, termed the Gaussian-sinc filter, to retain the complementary strengths of the Gaussian and sinc filters with no bandwidth or time expansion. The filter has lower sidelobes and is more localized than the sinc filter, enabling accurate I/O relation estimation and good data detection performance simultaneously. However, the Gaussian-sinc filter is not orthogonal to shifts by multiples of the delay and Doppler resolution, making it a sub-optimal choice.
In this paper, we design optimal DD pulse shaping filters using the Isotropic Orthogonal Transform Algorithm (IOTA)111while possible, we do not consider the alternate approach of Hermite pulse shapes [9, 10, 11] whose performance remains similar to IOTA pulse shapes [10], which has been extensively utilized in time-frequency multicarrier modulations (e.g., filterbank multicarrier) [12, 13, 10, 11]. We consider two maximally localized222satisfying Heisenberg’s Uncertainty Principle [13, 10, 14, 15] prototype pulse shapes – Gaussian and the prolate spheroidal wave function (PSWF) [14, 15] – and orthogonalize them to shifts by multiples of the delay and Doppler resolution via the IOTA procedure. In Section IV, we show that the proposed pulse shaping filters offer improved data detection and I/O relation estimation performance over the Gaussian-sinc filter from [8] since the latter is not orthogonal. Table I places our contributions in the context of prior work.
Notation: $x$ denotes a complex scalar, $\mathbf{x}$ denotes a vector with $n$th entry $\mathbf{x}[n]$, and $\mathbf{X}$ denotes a matrix with $(n,m)$th entry $\mathbf{X}[n,m]$. $(\cdot)^{\ast}$ denotes complex conjugate, $(\cdot)^{\top}$ denotes transpose, $(\cdot)^{\mathsf{H}}$ denotes complex conjugate transpose. Calligraphic font $\mathcal{X}$ denotes operators or sets, with usage clear from context. $|\cdot|$ denotes set cardinality or absolute value, with usage clear from context. $\mathbb{Z}$ denotes the set of integers and $\mathbb{Z}_{N}$ the set of integers modulo $N$. $\delta(\cdot)$ denotes the delta function, and $\mathbf{I}_{N}$ denotes the $N\times N$ identity matrix.
@@ II Overview of Zak-OTFS
We provide a brief overview of Zak-OTFS in this Section, and refer the interested reader to [3, 1, 2] for a more detailed description of Zak-OTFS.
The Zak-OTFS carrier waveform is a pulse in the delay-Doppler (DD) domain, formally a quasi-periodic localized function termed the DD pulsone. The DD pulsone is characterized by a delay period $\tau_{p}$ and a Doppler period $\nu_{p}$, with $\tau_{p}\nu_{p}=1$. The DD pulsone located at delay $\tau_{0}$ and Doppler $\nu_{0}$, where $0\leq\tau_{0}<\tau_{p},~0\leq\nu_{0}<\nu_{p}$, is given by [3, 1, 2]:
@@
$\mathbf{x}^{\mathbf{w}}_{{}_{\mathsf{DD}}}(\tau,\nu)$ $=\sum_{k_{0}=0}^{M-1}\sum_{l_{0}=0}^{N-1}\mathbf{X}[k_{0},l_{0}]\mathbf{p}_{(\nicefrac{{k_{0}}}{{B}},\nicefrac{{l_{0}}}{{T}})}^{\mathbf{w}}(\tau,\nu),$ (3)
where $\mathbf{X}$ denotes the $M\times N$ array of information symbols.
Note that from (2), we may equivalently express (3) as:
$\mathbf{x}^{\mathbf{w}}_{{}_{\mathsf{DD}}}(\tau,\nu)$ $=\mathbf{w}(\tau,\nu)*_{\sigma}\mathbf{x}_{{}_{\mathsf{DD}}}(\tau,\nu),$ (4)
in terms of an infinite time and bandwidth signal $\mathbf{x}_{{}_{\mathsf{DD}}}(\tau,\nu)$:
@@
$\mathbf{y}_{{}_{\mathsf{DD}}}(\tau,\nu)$ $=\mathbf{h}_{\mathrm{phy}}(\tau,\nu)*_{\sigma}\mathbf{x}^{\mathbf{w}}_{{}_{\mathsf{DD}}}(\tau,\nu)+\mathbf{n}_{{}_{\mathsf{DD}}}(\tau,\nu),$ (6)
where $\mathbf{n}_{{}_{\mathsf{DD}}}(\tau,\nu)$ denotes the additive noise at the receiver in the DD domain. The receiver performs matched filtering with the filter $\mathbf{\tilde{w}}(\tau,\nu)=e^{j2\pi\nu\tau}\mathbf{w}^{*}(-\tau,-\nu)$ to obtain [3, 1, 2]:
$\mathbf{y}^{\mathbf{\tilde{w}}}_{{}_{\mathsf{DD}}}(\tau,\nu)$ $=\mathbf{\tilde{w}}(\tau,\nu)*_{\sigma}\mathbf{y}_{{}_{\mathsf{DD}}}(\tau,\nu)$
$=\mathbf{h}_{\mathrm{eff}}(\tau,\nu)*_{\sigma}\mathbf{x}_{{}_{\mathsf{DD}}}(\tau,\nu)+\mathbf{n}^{\mathbf{\tilde{w}}}_{{}_{\mathsf{DD}}}(\tau,\nu),$ (7)
where the final step follows on substituting (4) in (6). In (II), $\mathbf{h}_{\mathsf{eff}}(\tau,\nu)$ denotes the effective channel555the effective channel approximates the physical channel when all paths are resolvable in delay with bandwidth $B$ and in Doppler with time $T$ that encompasses the effects of the physical scattering environment as well as pulse shaping and matched filtering [3, 1, 2]:
@@
and $\mathbf{n}^{\mathbf{\tilde{w}}}_{{}_{\mathsf{DD}}}(\tau,\nu)=\mathbf{\tilde{w}}(\tau,\nu)*_{\sigma}\mathbf{n}_{{}_{\mathsf{DD}}}(\tau,\nu)$ denotes the noise at the receiver after matched filtering.
Upon sampling on the information symbol lattice $\Lambda$, the system model in (II) reduces to [3, 1, 2]:
@@
where $\mathbf{y},\mathbf{x},\mathbf{n}$ are $MN\times 1$ vectors with $\mathbf{y}[k+lM]=\mathbf{y}^{\mathbf{\tilde{w}}}_{{}_{\mathsf{DD}}}[k,l]$, $\mathbf{x}[k+lM]=\mathbf{x}_{{}_{\mathsf{DD}}}[k,l]$, $\mathbf{n}[k+lM]=\mathbf{n}^{\mathbf{\tilde{w}}}_{{}_{\mathsf{DD}}}[k,l]$, and $\mathbf{H}$ is an $MN\times MN$ matrix with entries:
@@
Assuming the I/O relation (matrix $\mathbf{H}$) is known, the receiver detects the information symbols $\mathbf{x}$ from $\mathbf{y}$, e.g., via the minimum mean squared error (MMSE) estimator [3]. To acquire the I/O relation, a known pilot symbol is transmitted in one of three ways: in a frame separate from data [2], embedded into the same frame as data with appropriate guard regions [5, 6], or overlayed on the same frame as data via a mutually unbiased spreading filter [4]. The receiver estimates the sampled effective channel $\mathbf{h}_{\mathrm{eff}}[k,l]$ via the cross-ambiguity between the received symbols and the pilot symbols (see [4] for more details), which is subsequently used to estimate the matrix $\mathbf{H}$ via (II).
@@ III Pulse Shaping Filters & Their Design
III-A Pulse Shaping Filters in the Literature
Existing literature on Zak-OTFS [1, 2, 3, 4, 5, 6, 7, 8] has considered the following pulse shaping filters.
@@ III-A4 Gaussian-sinc [8]
III-B Proposed IOTA Pulse Shaping Design Framework
To overcome the drawbacks of existing pulse shaping filters from Section III-A (see Section I and Table I), we utilize the Isotropic Orthogonal Transform Algorithm (IOTA) [12, 13, 11, 10] framework to design DD pulse shaping filters. The key idea is to orthogonalize a maximally localized pulse shape to shifts of the lattice points $\Lambda$. The resulting pulse constitutes the closest orthogonal pulse shape to the initial maximally localized pulse.
@@ III-B Proposed IOTA Pulse Shaping Design Framework
Subsequently, we apply the IOTA procedure to the Gaussian and PSWF filters in two stages.
@@
$\tilde{\mathbf{G}}$ $=\mathbf{R}^{\nicefrac{{-1}}{{2}}}\mathbf{G}.$ (19)
The pulses $\tilde{\mathbf{G}}$ are orthogonal since the corresponding Gram matrix is an $MN\times MN$ identity matrix:
$\tilde{\mathbf{R}}$ $=\tilde{\mathbf{G}}\tilde{\mathbf{G}}^{\mathsf{H}}$
$=\mathbf{R}^{\nicefrac{{-1}}{{2}}}\mathbf{G}\mathbf{G}^{\mathsf{H}}\mathbf{R}^{\nicefrac{{-1}}{{2}}}$
$=\mathbf{I}_{MN}.$ (20)
@@ IV Numerical Results
(a) Pulse shaping filter magnitude vs normalized delay.
(b) Pulse shaping filter magnitude vs normalized Doppler.
(c) Energy spectral density (ESD) vs normalized frequency.
(d) Energy spectral density (ESD) vs normalized time.
@@
Relative power (dB) 0 -1 -9 -10 -15 -20
Figure 2: Uncoded 44-QAM data detection performance with perfect I/O relation knowledge at the receiver. The proposed IOTA pulses achieve ideal performance due to their orthogonality.
(a) Bit error rate (BER).
(b) Normalized mean squared error (NMSE).
Figure 3: Uncoded 44-QAM data detection performance with I/O relation estimation using a separate pilot frame with pilot SNR equal to data SNR. The simultaneous localization and orthogonality of the proposed IOTA pulses results in improved BER and I/O relation NMSE compared to the Gaussian-sinc filter [8], with no time or bandwidth expansion (unlike the RRC filter).
@@ IV-A Simulation Configuration
We conduct numerical simulations using a 3GPP-compliant $P=6$ path Vehicular-A (Veh-A) channel model [16] with significant mobility and delay spread, whose power-delay profile is shown in Table II. The Doppler of each path is simulated as $\nu_{i}=\nu_{\max}\cos(\theta_{i})$, with $\theta_{i}$ uniformly distributed in $\pi,\pi)$ and $\nu_{\max}=815$ Hz denoting the maximum channel Doppler spread. Note that our channel model is representative of real propagation environments since it considers fractional delay and Doppler shifts – the path delays in Table II being non-integer multiples of the delay resolution $\nicefrac{{1}}{{B}}$, and the Doppler shifts $\nu_{i}=\nu_{\max}\cos(\theta_{i})$ being non-integer multiples of the Doppler resolution $\nicefrac{{1}}{{T}}$. We simulate using the parameters: $M=17,N=19$, $\nu_{p}=30$ kHz and $\beta_{\tau}=\beta_{\nu}=0.6$ for the RRC filter. While simulations using larger frame sizes and wider bandwidths are possible, they are not pursued since we have shown in [17] that there is no difference in performance.

### v2

@@ Abstract
Zak-OTFS provides a framework for integrated sensing & communication (ISAC) in high delay and Doppler spread environments. Pulse shaping filter design enables joint optimization of sensing and communication performance. For sensing, a localized pulse shaping filter enables input-output (I/O) relation estimates close to the physical scattering channel. For communication, orthogonality of the pulse shape on the information lattice prevents inter-symbol interference, and no time and bandwidth expansion enables full spectral efficiency. A filter simultaneously meeting all three objectives is ideal for ISAC. Existing filter designs achieve two, but not all three objectives. In this work, we design pulse shaping filters meeting all three objectives via the Isotropic Orthogonal Transform Algorithm. The proposed filters have improved spectral efficiency, data detection and sensing performance over existing filter choices.
@@ I Introduction
Zak-OTFS (orthogonal time frequency space modulation) [15, 16, 17] is an emerging framework for integrated sensing and communication (ISAC) in doubly-selective environments with high delay and Doppler spreads. The Zak-OTFS carrier waveform is a pulse in the delay-Doppler (DD) domain, formally a quasi-periodic localized function termed pulsone, whose interaction with the doubly-selective channels does not fade [15, 16, 17]. This non-fading behavior of Zak-OTFS makes it well-suited for extreme mobility environments, such as bullet trains, where legacy time-frequency modulations, such as orthogonal frequency division multiplexing (OFDM), fade. Moreover, the response to a single Zak-OTFS carrier provides an image of the scattering environment, supporting sensing in radars [16] and advanced transportation networks [11].
In its original form, the DD pulsone occupies infinite time and bandwidth. For practical implementation, the DD pulsone is limited to a finite time interval $T$ and bandwidth $B$ via DD domain filtering with a pulse shaping filter [15, 16, 17, 22, 6, 7, 3, 2]. The choice of the pulse shaping filter is key to sensing and communication performance in interference-limited environments. A localized pulse shaping filter enables accurate input-output (I/O) relation estimation (sensing). Orthogonality of the pulse shape on the information lattice prevents inter-symbol and inter-slot interference (e.g., see [14]), and no time & bandwidth expansion enables full spectral efficiency. A filter simultaneously meeting all three objectives is ideal for ISAC.
Pulse shaping filters proposed in the Zak-OTFS literature [15, 16, 17, 22, 6, 7, 3, 2] meet two, but not all three objectives simultaneously. For instance, the sinc filter [15, 16, 17, 22, 6] is bandwidth & time limited and orthogonal on the information lattice, however, is not localized. Thus, it offers good data detection (communication) performance only given perfect I/O relation knowledge [15, 16, 17, 22, 6]. The root raised cosine (RRC) filter is localized and orthogonal, but is not bandwidth & time limited [15, 16, 17, 22, 6, 7, 3]. The Gaussian filter is maximally localized and bandwidth & time limited, but is not orthogonal [6], which degrades data detection performance. Recent work [2] utilizes the (normalized) product of the Gaussian and sinc filters, termed the Gaussian-sinc filter, which is more localized than the sinc filter and bandwidth & time limited, but is not orthogonal on the information lattice.
@@
In this paper, we design optimal DD pulse shaping filters using the Isotropic Orthogonal Transform Algorithm (IOTA)111while possible, we do not consider the alternate approach of Hermite pulse shapes [4, 19, 1] whose performance is similar to the Gaussian-sinc filter [8], by orthogonalizing the maximally localized222satisfying Heisenberg’s Uncertainty Principle [21, 19, 20, 9] Gaussian and the prolate spheroidal wave function (PSWF) [20, 9] pulse shapes. The IOTA procedure has been used to provide insight into characteristics of time-frequency modulations [10, 21, 19, 1]. Understanding the same characteristics of DD modulations is important and is novel to this paper. In Section IV, we show that the IOTA filters optimally balance the trade-off between localization and orthogonality without time & bandwidth expansion. Our results also show that while the Gaussian-sinc filter [2] is near-optimal, due to its non-orthogonality, additional spectral efficiency gains of $1\%$ are possible via orthogonalization using the IOTA approach. Table I places our contributions in the context of prior work.
@@ II Overview of Zak-OTFS
We provide a brief overview of Zak-OTFS in this Section, and refer the interested reader to [17, 15, 16] for more details.
The Zak-OTFS carrier waveform is a pulse in the delay-Doppler (DD) domain, formally a quasi-periodic localized function termed the DD pulsone. The DD pulsone is parameterized by a delay period $\tau_{p}$ and a Doppler period $\nu_{p}$, with $\tau_{p}\nu_{p}=1$. The DD pulsone located at delay $\tau_{0}$ and Doppler $\nu_{0}$, where $0\leq\tau_{0}<\tau_{p},~0\leq\nu_{0}<\nu_{p}$, is given by [17, 15, 16]:
@@
$\mathbf{x}^{\mathbf{w}}_{{}_{\mathsf{DD}}}(\tau,\nu)$ $=\sum_{k_{0}=0}^{M-1}\sum_{l_{0}=0}^{N-1}\mathbf{X}[k_{0},l_{0}]\mathbf{p}_{(\nicefrac{{k_{0}}}{{B}},\nicefrac{{l_{0}}}{{T}})}^{\mathbf{w}}(\tau,\nu)$
$=\mathbf{w}(\tau,\nu)*_{\sigma}\mathbf{x}_{{}_{\mathsf{DD}}}(\tau,\nu),$ (3)
where $\mathbf{X}$ is the $M\times N$ array of information symbols, and:
@@
$\mathbf{y}^{\mathbf{\tilde{w}}}_{{}_{\mathsf{DD}}}(\tau,\nu)$ $=\mathbf{h}_{\mathrm{eff}}(\tau,\nu)*_{\sigma}\mathbf{x}_{{}_{\mathsf{DD}}}(\tau,\nu)+\mathbf{n}^{\mathbf{\tilde{w}}}_{{}_{\mathsf{DD}}}(\tau,\nu),$ $=$ (5)
where $\mathbf{h}_{\mathsf{eff}}(\tau,\nu)$ denotes the effective channel555the effective channel approximates the physical channel when all paths are resolvable in delay with bandwidth $B$ and in Doppler with time $T$ that encompasses the effects of the physical scattering environment as well as pulse shaping and matched filtering [17, 15, 16]:
@@
and $\mathbf{n}^{\mathbf{\tilde{w}}}_{{}_{\mathsf{DD}}}(\tau,\nu)$ denotes receiver noise after matched filtering.
Upon sampling on the information symbol lattice $\Lambda$, the system model in (5) reduces to [17, 15, 16]:
@@
where $\mathbf{y},\mathbf{x},\mathbf{n}$ are $MN\times 1$ vectors with $\mathbf{y}[k+lM]=\mathbf{y}^{\mathbf{\tilde{w}}}_{{}_{\mathsf{DD}}}[k,l]$, $\mathbf{x}[k+lM]=\mathbf{x}_{{}_{\mathsf{DD}}}[k,l]$, $\mathbf{n}[k+lM]=\mathbf{n}^{\mathbf{\tilde{w}}}_{{}_{\mathsf{DD}}}[k,l]$, and $\mathbf{H}$ is the $MN\times MN$ I/O relation matrix with entries:
@@
Assuming the I/O relation matrix $\mathbf{H}$ is known, the receiver detects the information symbols $\mathbf{x}$ from $\mathbf{y}$, e.g., via the minimum mean squared error (MMSE) estimator [17]. To acquire the I/O relation, a known pilot symbol $\mathbf{x}_{\mathsf{p}}[k,l]$ is transmitted in one of three ways: in a frame separate from data [16], embedded in the data frame with appropriate guard regions [6, 7], or overlayed on the data frame via a mutually unbiased filter [22]. The receiver estimates the effective channel $\mathbf{h}_{\mathrm{eff}}[k,l]$ via the cross-ambiguity function [^7] Cross-ambiguity-based effective channel estimates are equivalent to range-Doppler maps in radar sensing. For more details, see [13, 17, 16, 22, 18]. between the received and transmitted pilot symbols [22]:
${\color[rgb]{0,0,0}\widehat{\mathbf{h}}_{\mathrm{eff}}[k,l]}$ ${\color[rgb]{0,0,0}=}{\color[rgb]{0,0,0}\frac{1}{MN}\sum_{k^{\prime}=0}^{M-1}\sum_{l^{\prime}=0}^{N-1}\mathbf{y}^{\mathbf{\tilde{w}}}_{{}_{\mathsf{DD}}}[k^{\prime},l^{\prime}]\mathbf{x}_{\mathsf{p}}^{*}[k^{\prime}-k,l^{\prime}-l]}$
$\qquad\qquad\qquad\quad{\color[rgb]{0,0,0}\times e^{-\frac{j2\pi}{MN}l(k^{\prime}-k)},}$ (10)
which is subsequently used to estimate the matrix $\mathbf{H}$ via (II).
@@ III Pulse Shaping Filters & Their Design
(a) Pulse shaping filter magnitude vs normalized delay.
(b) Energy spectral density (ESD) vs normalized frequency.
(c) Pulse shaping filter magnitude vs normalized Doppler.
(d) Energy spectral density (ESD) vs normalized time.
@@
III-A Standard Pulse Shaping Filters
The following pulse shaping filters have been considered in prior work on Zak-OTFS [15, 16, 17, 22, 6, 7, 3, 2].
@@ III-A4 Gaussian-sinc [2]
III-B IOTA Pulse Shaping Design Framework
To address the limitations of existing pulse shaping filters from Section III-A (see Section I and Table I), we utilize the Isotropic Orthogonal Transform Algorithm (IOTA) [10, 21, 1, 19] framework to design DD pulse shaping filters. The key idea is to orthogonalize a maximally localized pulse shape to shifts of the lattice points $\Lambda$. The resulting pulse constitutes the closest orthogonal pulse shape to the initially chosen pulse.
@@ III-B IOTA Pulse Shaping Design Framework
Subsequently, we apply the IOTA procedure in two stages.
@@
$\tilde{\mathbf{G}}$ $=\mathbf{R}^{\nicefrac{{-1}}{{2}}}\mathbf{G},$ (18)
with no inter-pulse interference on the lattice $\Lambda$ since:
$\tilde{\mathbf{R}}$ $=\tilde{\mathbf{G}}\tilde{\mathbf{G}}^{\mathsf{H}}=\mathbf{R}^{\nicefrac{{-1}}{{2}}}\mathbf{G}\mathbf{G}^{\mathsf{H}}\mathbf{R}^{\nicefrac{{-1}}{{2}}}=\mathbf{I}_{MN}.$ (19)
@@ IV Numerical Results
Relative power (dB) 0 1-1 9-9 10-10 15-15 20-20
@@ IV-A Simulation Configuration
We conduct numerical simulations using a 3GPP-compliant $P=6$ path Vehicular-A (Veh-A) channel model [5] with significant mobility and delay spread. The channel gain of each path is sampled from the power-delay profile in Table II with random phases, and the Doppler of each path is simulated as $\nu_{i}=\nu_{\max}\cos(\theta_{i})$, with $\theta_{i}$ uniformly distributed in $\pi,\pi)$ and $\nu_{\max}$ denoting the maximum channel Doppler spread [^8] Note that our channel model is representative of real propagation environments with fractional delay and Doppler shifts – the path delays in Table II are non-integer multiples of the delay resolution $\nicefrac{{1}}{{B}}$, and the Doppler shifts $\nu_{i}=\nu_{\max}\cos(\theta_{i})$ are non-integer multiples of the Doppler resolution $\nicefrac{{1}}{{T}}$.. We simulate using the parameters: $M=17,N=19$, $\nu_{\max}=815$ Hz999at a carrier frequency of $3$ GHz, this corresponds to a highly dynamic scenario with maximum speed of $81.5$ m/s, e.g., a high-speed train, $\nu_{p}=30$ kHz and $\beta_{\tau}=\beta_{\nu}=0.6$ for the RRC filter. We do not pursue simulations with larger frame sizes and wider bandwidths here, since we have shown there is no difference in performance in [12].
Figure 2: Normalized mean squared error (NMSE) for estimating the DD effective channel $\mathbf{h}_{\mathrm{eff}}[k,l]$ (sensing) via (II).
(a) Bit error rate (BER) with perfect I/O relation knowledge.
(b) BER with estimated I/O relation.
Figure 3: Uncoded 44-QAM data detection performance with perfect and estimated I/O relation. The simultaneous localization, orthogonality and improved I/O relation estimation with the proposed IOTA pulses results in near-ideal BER performance.